A particle is moving in the +X direction with its velocity  V=ax^1/2 [a is constant] .Assume that at t=0,the particle was located at X=0 then the value of time dependence of velocity

Grade:11

1 Answers

deepak
7 years ago
v=ax^1/2
then, 
dx/dt=v=a(x^1/2)
dx/(x^1/2)=a(dt)
on intergrating , we get ,
2(x^1/2)=at+c
since at t=0 => x=0 => c=0
therefore,
2(x^1/2)=at        
subs, in the first eqn,
v=a(at)/2